Respected sir, Qn. Given,6 M HCN in 1ltr solotion of water , Find density in g/cc. If DENSITY is expressed a D how 1000D has 6 moles HCN?Pls explain how 1000 D we get? Thnx

Respected sir,
Qn. Given,6 M HCN in 1ltr solotion of water , Find density in g/cc.
If DENSITY is  expressed a D how 1000D has 6 moles HCN?Pls explain how 1000 D we get?
 
Thnx

Grade:11

1 Answers

Sahani Kumar
5 years ago
Molecular mass of HCN = 27 gram
As 1 mole of HCN = 27 gram
So 6 mole of HCN will be = (6 * 27) = 162 gram
6 molar solution  of HCN means that 
6 mole of HCN are present in 1 Litre solution( or 1000 cc).
As we know that 6 mole of HCN = 162 gram 
so we can say that
1000 cc solution contain 162 gram of HCN 
So 1 cc solution will contain= (162/1000)g/cc
                                                  = 0.162g/cc
Regards 
Askiitians Member
If you like this answer please encourage us by approving this answer.